OpenStudy (anonymous):

help me with this fraction problem please.... 3/7 - 4/m

OpenStudy (anonymous):

what do we need to do?

OpenStudy (anonymous):

is it \(\frac{3}{7}-\frac{4}{m}\) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and you have to put over one denominator, then just like if m was a number you have to write with the common denominator of \(7m\) and write \[\frac{3\times m-4\times 7}{7m}\]

OpenStudy (anonymous):

clean it up as \[\frac{3m-28}{7m}\]

OpenStudy (anonymous):

ok so is this the answer?